# Simultaneous Khintchine theorem on manifolds in positive characteristics: convergence case

### v1

@@ 1.2. Main Results
Let $d+m=n\geq 2$, $\bm{\Theta}\in\mathcal{K}^{n}$ and $\psi:(0,\infty)\rightarrow(0,1)$ be an approximating function. Also let $\mathscr{M}\subset\mathcal{K}^{n}$ be an analytic submanifold satisfying (I), (II), and (IV). Suppose that $\sigma>0$ is such that
$\sum_{t\geq 1}\left(\frac{\psi(q^{t})}{q^{t}}\right)^{\sigma+m}q^{(n+1)t}<\infty,$ (1.2)
and
$\sum_{t\geq 1}\left(\frac{\psi(q^{t})}{q^{t}}\right)^{\frac{\sigma-d}{2}}<\infty.$ (1.3)
Then,
$\mathcal{H}^{\sigma}\left(\mathbf{f}^{-1}(\mathcal{S}_{n}^{\bm{\Theta}}(\psi))\right)=0.$
Given integers $t,s>0$ and $\Delta\subseteq\mathcal{K}^{n}$, we define the following set, which counts the number of rational points $\frac{{\mathbf{P}}}{Q}$ that are within a distance $q^{-(s+t)}$ of $\mathbf{f}(\Delta\cap U)$,
$\mathcal{R}_{\bm{\Theta}}(\Delta,s,t):=$ $\left\{(\mathbf{P},Q)\in\mathcal{R}^{n}\times(\mathcal{R}\setminus\{0\}):|Q|=q^{t}\ \text{and}\ \inf_{x\in\Delta\cap U}\left\|\mathbf{f}(\mathbf{x})-\frac{{\mathbf{P}}+\bm{\Theta}}{Q}\right\|<q^{-(s+t)}\right\},$
and define $\mathcal{N}_{\bm{\Theta}}(\Delta;s,t)=\#\mathcal{R}_{\bm{\Theta}}(\Delta,s,t)$.
To prove Theorems 1.2 and 1.3, we divide the manifold into two parts: we count the number of rational points near the manifold in the generic part ($U\setminus\mathfrak{M}(s,t)$) and show that the measure of the special part ($\mathfrak{M}(s,t)$) of the manifold is relatively small. For the definition of $\mathfrak{M}(s,t)$, we refer the reader to subsection 4.2. In fact, we prove the following proposition.
Proposition 1.
Let $U\subseteq\mathcal{K}^{d}$ be an open set, and let $\mathbf{f}:U\to\mathcal{K}^{n}$ be a map satisfying (I), (II), and (IV). Then, for any $s,t>0$, there exists a set $\mathfrak{M}(s,t)\subseteq U$ that can be written as the union of disjoint balls in $U$ of radius $q^{-\frac{s+t}{2}}$. For every $\mathbf{x}_{0}\in U$, there exists a ball $B_{0}$ centered at $\mathbf{x}_{0}$ and constants $C,\alpha$ (depending on $B_{0},\mathbf{f}$) such that for every sufficiently large $t>0$, we have
$\mathcal{L}_{d}(\mathfrak{M}(s,t)\cap B_{0})\leq C\max\left\{q^{-\alpha t},q^{\alpha\left(\left(\frac{2n-1}{2(n+1)}\right)s-\frac{3}{2(n+1)}t\right)}\right\}\mathcal{L}_{d}(B_{0}),$
and for every ball $B\subseteq U$ and for every sufficiently large $t$, we have
$\mathcal{N}_{\bm{\Theta}}(B\setminus\mathfrak{M}(s,t),s,t)\ll_{n,\mathbf{f}}q^{t}\frac{q^{dt}}{q^{ms}}\mathcal{L}_{d}(B),$
where $A\ll_{n,\mathbf{f}}B$ means there exists a constant $c>0$ depending on $n,\mathbf{f}$ such that $A\leq cB.$
@@ 1.3. Counting Heuristics
We now motivate the heuristic count leading to Proposition 1. Let $Q\in\mathcal{R}$ be a polynomial with $\text{deg}\ Q=t$. Then, the number of such polynomials $Q$ is
$\{Q\in\mathcal{R}:\deg Q=t\}=(q-1)q^{t-1}\asymp q^{t}.$
For a fixed denominator $Q$ and a fixed $i\in\{1,\dots,n\}$, the number of $P_{i}\in\mathcal{R}$ with $\deg(P_{i})\leq t$ is $q^{t+1}$. Therefore,
@@
Heuristically, this is simplified as follows.
$q^{n(t+1)}=q^{n}\cdot q^{nt}\ll q^{nt},$
since $q^{n}$ is a constant independent of $t$. Hence, the total number of rational points of the form $\frac{{\mathbf{P}}}{Q}\in\mathbb{F}_{q}(T)^{n}$ with $\deg Q=t$, is
$\ll q^{t}\cdot q^{nt}=q^{(n+1)t}.$
Hence, outside a set of small measure, that is the set $\mathfrak{M}(s,t)$ from Proposition 1, the number of rational points $\frac{{\mathbf{P}}}{Q}$ that are within distance $q^{-(s+t)}$ of $\mathbf{f}(\Delta\cap U)$ is
$\ll(q^{-(s+t)})^{m}q^{(n+1)t}=q^{-ms}q^{(d+1)t}.$
@@ 1.4. Constraints and Differences in the Positive Characteristic Setting
Although our methodology largely follows the general framework introduced in [BY23], the function field setting poses new difficulties which require a fresh perspective. In the real case, the simultaneous form of Khintchine’s theorem established in [BY23] relies on diagonal flows whose entries are fractional powers of the form $e^{t/\ell}$. The natural analogue in the function field context would involve powers of $T^{\frac{1}{\ell}}$, which do not belong to the base field $\mathcal{K}$, but rather to its extension $\mathcal{K}(T^{\frac{1}{\ell}})$.
To adapt the approach of [BY23], we therefore work over the extended field $\mathcal{K}(T^{\frac{1}{\ell}})$. This, however, introduces additional challenges. In particular, the polynomial ring $\mathcal{R}$ is no longer a lattice in the extension $\mathcal{K}(T^{\frac{1}{\ell}})$, but merely a discrete subgroup. As a result, the classical tools of the geometry of numbers are not directly applicable. To overcome this obstacle, we establish results in the geometry of numbers for specific discrete subgroups, which are not necessarily lattices (see section˜3). Moreover, in the appendix section˜6, we establish a result concerning the first successive minima of such discrete subgroups over $\mathcal{K}(T^{\frac{1}{\ell}})$, which we believe to be of independent interest.
@@ 1.5. Structure of the Paper
We begin in section˜2 by recalling the definitions of analytic and $C^{k}$ functions in the ultrametric setting. In section˜3, we establish several auxiliary results in the geometry of numbers over function fields. Proposition 1 is proved in §4 by deriving a measure estimate for the special part of the manifold (see section˜4.3) and a counting estimate for the generic part (see section˜4.2). Finally, Proposition 1 serves as the key ingredient in the proofs of Theorems 1.2 and 1.3, which are presented in section˜5. We conclude the paper with an additional result on the geometry of numbers over function fields, proved in §6.
@@ Acknowledgements
The authors express their sincere gratitude to the University of York for hosting the workshop "Diophantine Approximation and Related Fields," during which a part of this work was done. The authors thank Prof. Anish Ghosh for his continuous encouragement. The fourth-named author would like to thank Prof. Ravi Raghunathan, Prof. Keshav Aggarwal, and Prof. Dipendra Prasad for their interest in this project. The fourth-named author is supported by the Prime Minister’s Research Fellowship (ID: 1302639) during the course of the work.
@@ 2. Ultrametric Calculus
In this section, we recall the concept of ultrametric $C^{k}$ functions from [Sch84]. Let $U$ be an open subset of $\mathcal{K}$, and $g:U\subseteq\mathcal{K}\to\mathcal{K}$ be a function. The first-order difference quotient of $g$, denoted by $\Phi^{1}g$, is defined as
$\Phi^{1}g(x,y):=\frac{g(x)-g(y)}{x-y},\quad{for}\,\,(x,y)\in\nabla^{2}U,$
where
$\nabla^{2}U:=\{(x,y)\in U\times U\mid x\neq y\}.$
We say that $g$ is $C^{1}$ at a point $a\in U$ if the limit
$\lim_{(x,y)\to(a,a)}\Phi^{1}g(x,y)$
exists and $g$ is called $C^{1}$ on $U$ if it is $C^{1}$ at every point of $U$. To define a $C^{k}$ function, consider
$\nabla^{k}U:=\{(x_{1},\dots,x_{k})\in U^{k}\mid x_{i}\neq x_{j}\text{ for }i\neq j\},$
and define the $k$-th order difference quotient $\Phi^{k}g:\nabla^{k+1}U\rightarrow\mathcal{K}$ inductively:
$\Phi^{0}g:=g,\quad\Phi^{k}g(x_{1},\dots,x_{k+1}):=\frac{\Phi^{k-1}g(x_{1},x_{3},\dots,x_{k+1})-\Phi^{k-1}g(x_{2},x_{3},\dots,x_{k+1})}{x_{1}-x_{2}}.$
One can easily see that $\Phi^{k}g$ is a symmetric function of its $k+1$ variables. $g$ is said to be $C^{k}$ at a point $a\in U$ if the following limit exists:
$\lim_{(x_{1},\dots,x_{k+1})\to(a,\dots,a)}\Phi^{k}g(x_{1},\dots,x_{k+1})$
and $g$ is said to be $C^{k}$ on $U$, i.e., $g\in C^{k}(U)$ if $g$ is $C^{k}$ at every point of $U$. This is equivalent to $\Phi^{k}g$ admitting an extension $\bar{\Phi}^{k}g:U^{k+1}\rightarrow{\mathcal{K}}$, which, if it exists, is unique. If $g\in C^{k}(U)$, all derivatives up to order $k$ exist and the $k$-th derivative of $g$ is given by:
$g^{(k)}(x)=k!\,\bar{\Phi}^{k}g(x,\dots,x).$
Now let $g:U_{1}\times\dots\times U_{d}\subseteq\mathcal{K}^{d}\rightarrow\mathcal{K}$ be a $\mathcal{K}$-valued function of several variables, where $U_{i}\subseteq\mathcal{K}$ are open subsets for $i=1,\dots,d$. Let $\Phi_{i}^{k}g$ denote the $k$-th order difference quotient with respect to the $i$-th variable. For a multi-index $\beta=(i_{1},\dots,i_{d})$, we define:
$\Phi_{\beta}g:=\Phi_{1}^{i_{1}}\circ\cdots\circ\Phi_{d}^{i_{d}}g.$
The domain of $\Phi_{\beta}g$ is $\nabla^{i_{1}+1}U_{1}\times\cdots\times\nabla^{i_{d}+1}U_{d}.$ We say $g\in C^{k}(U_{1}\times\cdots\times U_{d})$ if for all multiindex $\beta=(i_{1},\dots,i_{d})$ with $|\beta|:=\sum_{j=1}^{d}i_{j}\leq k$, the difference quotient $\Phi_{\beta}g$ extends continuously to
$\bar{\Phi}_{\beta}g:U_{1}^{i_{1}+1}\times\cdots\times U_{d}^{i_{d}+1}\to\mathcal{K}.$
As in the one-variable case, for $g\in C^{k}(U_{1}\times\dots\times U_{d})$, the mixed partial derivatives
$\partial_{\beta}g:=\partial_{1}^{i_{1}}\circ\cdots\circ\partial_{d}^{i_{d}}g$
exist and are continuous for all $|\beta|\leq k$. They also satisfy:
@@
where each variable $x_{j}$ appears $i_{j}+1$ times, and $\beta!:=\prod_{j=1}^{d}i_{j}!$. Now we define an analytic function in $\mathcal{K}.$
@@
Let $U$ be an open subset of $\mathcal{K}$. A function $g:U\subseteq\mathcal{K}\to\mathcal{K}$ is called analytic on $U$ if for every point $x_{0}\in U$, there exists an open ball
$B(x_{0},r)=\{x\in\mathcal{K}:|x-x_{0}|<r\}\subseteq U$
and a power series
$g(x)=\sum_{n=0}^{\infty}a_{n}(x-x_{0})^{n}$
with coefficients $a_{n}\in\mathcal{K}$, such that this series converges for all $x\in B(x_{0},r)$, and equals $g(x)$ on that ball. In other words, $g$ is analytic on $U$ if it is locally given by a convergent power series at every point in $U$.
Given an analytic map $g=(g_{1},\dots,g_{n}):U\subseteq\mathcal{K}^{d}\to\mathcal{K}^{n}$, by $\nabla g(\mathbf{x})$ we denote the $d\times n$ matrix whose $(i,j)$-th entry is $\partial_{j}g_{i}(x)$. We also recall the following second-order Taylor expansion formula for analytic functions (see [BDG24, §11-Appendix] for details).
@@
Let $U$ be an open subset of $\mathcal{K}^{d}$ and $\mathbf{x},\mathbf{x}^{\prime}\in U$. Also, let $g:U\to\mathcal{K}^{n}$ be an analytic function. Then, we have
$g(\mathbf{x}+\mathbf{x}^{\prime})=g(\mathbf{x})+\sum_{i=1}^{d}x_{i}^{\prime}\bar{\Phi}_{e_{i}}g(\cdot)+\sum_{\beta=(i_{1},\ldots,i_{d}),|\beta|=2}\bar{\Phi}_{\beta}g(\cdot)\prod_{i=1}^{d}(x_{i}^{\prime})^{i_{j}},$
where $e_{i}$ is the multiindex whose $i$-th coordinate is $1$ and all other coordinates are zero, and the arguments of $\bar{\Phi}_{e_{i}}g(\cdot)$ and $\bar{\Phi}_{\beta}g(\cdot)$ are some of the components of $\mathbf{x}$ and $\mathbf{x}^{\prime}$.
In particular, in our context, applying the above formula for $f_{i}$’s and using the bounds of the first and second difference quotients of $f_{i}$ (see (IV)), we have
$\|f_{i}(\mathbf{x}+\mathbf{x}^{\prime})-f_{i}(\mathbf{x})\|\leq\max\{M\|\mathbf{x}^{\prime}\|,M\|\mathbf{x}^{\prime}\|^{2}\}\,\,\,\,\text{for}\,\,i=1,\dots,m.$ (2.2)
We conclude this section with the following version of Besicovitch’s covering theorem, which will be used in section˜4.
@@
[KT07, Example 2.1] Let $A$ be a bounded subset of $\mathcal{K}^{d}$. Then any covering by balls of $A$ has a countable subcover consisting of mutually disjoint balls.

### v2

@@ 1.2. Main Results
Let $d+m=n\geq 2$, $\bm{\Theta}\in\mathcal{K}^{n}$, and $\psi:\{q^{r}:r\in\mathbb{Z}_{\geq 0}\}\longrightarrow\{q^{r}:r\in\mathbb{Z}\}$ be an approximating function. Suppose that $\mathscr{M}\subseteqq\mathcal{K}^{n}$ is an analytic submanifold satisfying ($\mathfrak{A}$.1)-($\mathfrak{A}$.4). Assume that $\sigma>0$ and $\alpha>0$ are such that the following pair of series is convergent:
$\sum_{Q\in\mathbb{F}_{q}[T]\setminus\{0\}}\left(\frac{\psi(|Q|)}{|Q|}\right)^{\sigma+m}|Q|^{n}\text{ and }\sum_{Q\in\mathbb{F}_{q}[T]\setminus\{0\}}\left(\frac{\psi(|Q|)}{|Q|}\right)^{\frac{\sigma-d}{2}}\left(\psi(|Q|)^{-\frac{n\alpha}{n+1}}|Q|^{-\frac{3\alpha}{2(n+1)}-1}\right).$ (1.5)
Then, $\mathcal{H}^{\sigma}\left(\mathbf{f}^{-1}(\mathcal{S}_{n}^{\bm{\Theta}}(\psi))\right)=0$.
We prove both the above theorems in §5. Similar to [BY23], the proof strategy is to divide the manifold into two parts: we count the number of rational points near the manifold in the generic part $U\setminus\mathfrak{M}(s,t)$ and show that the measure of the special part $\mathfrak{M}(s,t)$ of the manifold is relatively small. We refer the reader to §4.2 for the definition of $\mathfrak{M}(s,t)$.
Given integers $t,s>0$ and $\Delta\subseteqq\mathcal{K}^{n}$, we define the following set, which counts the number of rational points $\frac{{\mathbf{P}}}{Q}$ that are within a distance $q^{-(s+t)}$ of $\mathbf{f}(\Delta\cap U)$:
$\mathcal{R}_{\bm{\Theta}}(\Delta,s,t){\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}$ $\left\{(\mathbf{P},Q)\in\mathcal{R}^{n}\times(\mathcal{R}\setminus\{0\}):|Q|=q^{t}\text{ and }\inf_{\mathbf{x}\in\Delta\cap U}\left\|\mathbf{f}(\mathbf{x})-\frac{{\mathbf{P}}+\bm{\Theta}}{Q}\right\|<q^{-(s+t)}\right\},$
and $\mathcal{N}_{\bm{\Theta}}(\Delta;s,t){\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\#\mathcal{R}_{\bm{\Theta}}(\Delta,s,t)$. We prove the following proposition.
Proposition 1.1.
Let $U\subseteqq\mathcal{K}^{d}$ be an open set, and let $\mathbf{f}:U\to\mathcal{K}^{n}$ be a map satisfying ($\mathfrak{A}$.1)-($\mathfrak{A}$.4), and $1<\kappa<\frac{3}{2}$. Then there exists a family $\left\{\mathfrak{M}(s,t):s,t\in\mathbb{N}\right\}$ of subsets of $U$ satisfying the following:
  1. ($\mathfrak{S}$.1)
    $\mathfrak{M}(s,t)$ can be expressed as the union of disjoint balls in $U$ with radius $q^{-\frac{s+t}{2}}$;
  2. ($\mathfrak{S}$.2)
    for every $\mathbf{x}_{0}\in U$, there exists a ball $B_{0}$ centered at $\mathbf{x}_{0}$ and positive constants $C$ and $\alpha$, depending on $B_{0}$ and $\mathbf{f}$, such that
    $\mathcal{L}_{d}(\mathfrak{M}(s,t)\cap B_{0})\leq Cq^{\frac{\alpha}{n+1}\left(ns-\frac{3t}{2}\right)}\mathcal{L}_{d}(B_{0}),\text{ whenever }t\gg 1\text{ and }s\leq\frac{\kappa t}{n};\text{ and }$
  3. ($\mathfrak{S}$.3)
    for every ball $B\subseteqq U$, the following holds for all sufficiently large $t$:
    $\forall s\in\mathbb{N},\penalty 10000\ \mathcal{N}_{\bm{\Theta}}(B\setminus\mathfrak{M}(s,t),s,t)\ll_{n,\mathbf{f}}q^{(d+1)t-ms}\mathcal{L}_{d}(B),$
    where $A\ll_{n,\mathbf{f}}B$ means that there exists a constant $c>0$ depending on $n,\mathbf{f}$ such that $A\leq cB.$
@@ 1.3. Counting Heuristics
We now motivate the heuristic count leading to Proposition 1.1. Let $Q\in\mathcal{R}$ with $\deg Q=t$. Then
$\{Q\in\mathcal{R}:\deg Q=t\}=(q-1)q^{t}\asymp q^{t+1}.$
For a fixed denominator $Q$ and a fixed $i\in\{1,\dots,n\}$, the number of polynomials $P_{i}\in\mathcal{R}$ with $\deg P_{i}\leq t$ is $q^{t+1}.$ Therefore,
@@
Hence, for denominators of degree exactly $t$, the total number of rational points of the form $\frac{\mathbf{P}}{Q}\in\mathbb{F}_{q}(T)^{n}$ is
$(q-1)q^{t}\cdot q^{n(t+1)}=(q-1)q^{t+n(t+1)}=(q-1)q^{(n+1)t+n}\asymp q^{(n+1)t}.$
Outside the exceptional set $\mathfrak{M}(s,t)$ from Proposition 1.1, the number of rational points $\frac{\mathbf{p}}{Q}$ lying within a distance $q^{-(s+t)}$ of $\mathbf{f}(\Delta\cap U)$ is heuristically bounded by
$(q^{-(s+t)})^{m}\,q^{(n+1)t}=q^{-ms}q^{(d+1)t}.$
@@ 1.4. Constraints and Differences in the Positive Characteristic Setting
While our approach is similar to that of [BY23] and [BD25], the function field setting poses new challenges that require a fresh perspective, given the non-monotonicity of the approximating function $\psi$. In the real case, one typically divides $1,\infty)$ into the intervals $\left[e^{t-1},e^{t}\right]$, where $t\in\mathbb{N}$, so that every $\psi$-approximable vector $\mathbf{y}\in\mathbb{R}^{n}$ satisfies
$\left\|\mathbf{y}-\frac{\mathbf{p}}{q}\right\|_{\infty}<\frac{\psi(e^{t-1})}{e^{t-1}},\text{ with }\mathbf{p}\in\mathbb{Z}^{n},q\in\mathbb{N},\text{ and }e^{t-1}\leq q<e^{t},$
for infinitely many $t\in\mathbb{N}$. This reduction, which paves the way for the application of the Borel-Cantelli lemma, requires some form of monotonicity. In our scenario, the lack of any monotonicity assumption in the function field setting forces us to perform heightwise counting, i.e., we count polynomials $Q$ with $|Q|=q^{t}$ (see Lemma 3.7 for details). To this end, it is crucial that the main term of the upper bound for rational point counting coincides with the heuristic bound mentioned in §1.3. This is really difficult to avoid overcounting if the entries of diagonal flow remain in $\mathcal{K}$, particularly with integral powers of $T$. Moreover, another potential issue with working with integral powers is that sensible counting can only be performed for $t$s that belong to some proper infinite subset of $\mathbb{N}$. This limitation results in a smaller $\limsup$ set than what is required.
To circumvent this, we therefore work over the extended field $\mathcal{K}\left(T^{\frac{1}{\ell}}\right)$. This, however, introduces additional challenges; for instance, the polynomial ring $\mathcal{R}$ is no longer a lattice in the extension $\mathcal{K}\left(T^{\frac{1}{\ell}}\right)$, but merely a discrete subgroup. As a result, the classical tools of the geometry of numbers are not directly applicable. To overcome this obstacle, we establish results in the geometry of numbers for specific discrete subgroups that are not necessarily lattices (see section˜3). Moreover, in the appendix section˜6, we establish a result concerning the first successive minima of such discrete subgroups over $\mathcal{K}\left(T^{\frac{1}{\ell}}\right)$, which we believe to be of independent interest.
@@ 1.5. Structure of the Paper
We begin in section˜2 by recalling the definitions of analytic and $C^{k}$ functions in the ultrametric setting. In section˜3, we establish several auxiliary results in the geometry of numbers over function fields. Proposition 1.1 is proved in §4 by deriving a measure estimate for the special part of the manifold (see section˜4.3) and a counting estimate for the generic part (see section˜4.2). Finally, Proposition 1.1 serves as the key ingredient in the proofs of Theorems 1.2 and 1.4, which are presented in section˜5. We conclude the paper with an additional result on the geometry of numbers over function fields, proved in §6.
@@ Acknowledgements
The authors express their sincere gratitude to the University of York for hosting the workshop “Diophantine Approximation and Related Fields,” during which part of this work was completed. The authors thank Anish Ghosh for his continuous encouragement. The authors would like to thank Shreyasi Datta and Subhajit Jana for many helpful discussions and insightful comments. The fourth-named author is supported by the Prime Minister’s Research Fellowship (ID: 1302639) during the course of the work.
@@ 2. Ultrametric Calculus
In this section, we recall the concept of ultrametric $C^{k}$ functions from [Sch84, Chapter 2]. Let $U$ be an open subset of $\mathcal{K}$, and $g:U\longrightarrow\mathcal{K}$. The first-order difference quotient of $g$, denoted by $\Phi^{1}g$, is defined as
$\Phi^{1}g(x,y){\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\frac{g(x)-g(y)}{x-y},\penalty 10000\ \forall(x,y)\in\nabla^{2}U{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\{(x,y)\in U\times U\mid x\neq y\}.$
We say that $g$ is $C^{1}$ at a point $a\in U$ if $\lim_{(x,y)\to(a,a)}\Phi^{1}g(x,y)$ exists, and $g$ is called $C^{1}$ on $U$ if it is $C^{1}$ at every point of $U$. Note that $g\in C^{1}(U)$ if and only if $\Phi^{1}g$ can be uniquely extended to a continuous function $\bar{\Phi}^{1}g$ on $U\times U$. The concept of a $C^{k}$ function can be defined inductively. Indeed, given any $k\in\mathbb{N}$, let $\nabla^{k}U{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\{(x_{1},\dots,x_{k})\in U^{k}\mid x_{i}\neq x_{j}\text{ for }i\neq j\}$. Define the $k$-th order difference quotient $\Phi^{k}g:\nabla^{k+1}U\longrightarrow\mathcal{K}$ inductively:
$\Phi^{0}g{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}g,\quad\Phi^{k}g(x_{1},\dots,x_{k+1}){\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\frac{\Phi^{k-1}g(x_{1},x_{3},\dots,x_{k+1})-\Phi^{k-1}g(x_{2},x_{3},\dots,x_{k+1})}{x_{1}-x_{2}}.$
One can easily see that $\Phi^{k}g$ is a symmetric function of its $k+1$ variables. The function $g$ is said to be $C^{k}$ at a point $a\in U$ if the following limit exists:
$\lim_{(x_{1},\dots,x_{k+1})\to(a,\dots,a)}\Phi^{k}g(x_{1},\dots,x_{k+1}),$
and $g$ is said to be $C^{k}$ on $U$, which is denoted by $g\in C^{k}(U)$, if $g$ is $C^{k}$ at every point of $U$. In view of [Sch84, Theorem 29.9], this is equivalent to $\Phi^{k}g$ admitting a continuous extension $\bar{\Phi}^{k}g$ to $U^{k+1}$, which must be unique as $\nabla^{k}U$ is dense in $U^{k}$. It follows that if $g\in C^{k}(U)$, all derivatives up to order $k$ exist and are continuous. In fact, one has
$g^{(k)}(\mathbf{x})=k!\,\bar{\Phi}^{k}g(x,\dots,x),\penalty 10000\ \forall x\in U.$
Now let $g:U_{1}\times\dots\times U_{d}\subseteqq\mathcal{K}^{d}\longrightarrow\mathcal{K}$, where $U_{i}\subseteqq\mathcal{K}$ are open subsets for $i=1,\dots,d$. Denote by $\Phi_{i}^{k}g$ the $k$-th order difference quotient with respect to the $i$-th variable. For a multi-index $\beta=(i_{1},\dots,i_{d})$, we define the following:
$\Phi_{\beta}g{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\Phi_{1}^{i_{1}}\circ\cdots\circ\Phi_{d}^{i_{d}}g.$
The domain of $\Phi_{\beta}g$ is $\nabla^{i_{1}+1}U_{1}\times\cdots\times\nabla^{i_{d}+1}U_{d}.$ We say $g\in C^{k}(U_{1}\times\cdots\times U_{d})$ if for all multi-indices $\beta=(i_{1},\dots,i_{d})$ with $|\beta|{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\sum_{j=1}^{d}i_{j}\leq k$, the difference quotient $\Phi_{\beta}g$ extends continuously to a function $\bar{\Phi}_{\beta}g$ on $U_{1}^{i_{1}+1}\times\cdots\times U_{d}^{i_{d}+1}$. As in the one-variable case, for $g\in C^{k}(U_{1}\times\dots\times U_{d})$, the mixed partial derivatives $\partial_{\beta}g{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\partial_{1}^{i_{1}}\circ\cdots\circ\partial_{d}^{i_{d}}g$ exist and are continuous for all $|\beta|\leq k$. Indeed, it is clear that
@@
where each variable $x_{j}$ appears $i_{j}+1$ times and $\beta!{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\prod_{j=1}^{d}i_{j}!$.
@@
Let $U$ be an open subset of $\mathcal{K}^{d}$. A function $g:U\subseteqq\mathcal{K}^{d}\to\mathcal{K}$ is called analytic on $U$ if for every point $\mathbf{x}_{0}\in U$, there exists an open ball $B(\mathbf{x}_{0},r)\subseteqq U$ and a power series $\sum_{\beta\in\mathbb{N}_{0}^{d}}a_{\beta}(\mathbf{x}-\mathbf{x}_{0})^{\beta}$, with coefficients $a_{\beta}\in\mathcal{K}$ and $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$, such that
[…]
where, given any multi-index $\beta=(i_{1},\dots,i_{d})\in\mathbb{N}_{0}^{d}$ and $\mathbf{y}=(y_{1},\dots,y_{d})\in\mathcal{K}^{d}$, $\mathbf{y}^{\beta}{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}y_{1}^{i_{1}}\dots y_{d}^{i_{d}}$. In other words, $g$ is analytic on $U$ if it is locally given by a convergent power series about every point in $U$.
Note that any analytic function is of class $C^{\infty}$. Given an analytic map $\mathbf{g}:U\subseteqq\mathcal{K}^{d}\longrightarrow\mathcal{K}^{n}$ with coordinates $g_{1},\dots,g_{n}$, by $\nabla\mathbf{g}(\mathbf{x})$ we denote the $d\times n$ matrix whose $(i,j)$-th entry is $\partial_{j}g_{i}(\mathbf{x})$. We also recall the following second-order Taylor’s formula for analytic functions ([BDG24, §11-Appendix]).
@@
Let $U$ be an open subset of $\mathcal{K}^{d}$, and $\mathbf{x}\in U$ and $\mathbf{h}\in\mathcal{K}^{d}$ are such that $\mathbf{x}+\mathbf{h}\in U$. Then for any analytic $g:U\longrightarrow\mathcal{K}$,
$g(\mathbf{x}+\mathbf{h})=g(\mathbf{x})+\sum_{i=1}^{d}h_{i}\bar{\Phi}_{e_{i}}g(\cdot)+\sum_{\beta=(i_{1},\ldots,i_{d}),|\beta|=2}\bar{\Phi}_{\beta}g(\cdot)\prod_{j=1}^{d}(h_{j})^{i_{j}},$ (2.2)
where $e_{i}$ is the multi-index whose $i$-th coordinate is $1$ and all other coordinates are zero, and the arguments of $\bar{\Phi}_{e_{i}}g(\cdot)$ and $\bar{\Phi}_{\beta}g(\cdot)$ are some of the components of $\mathbf{x}$ and $\mathbf{h}$.
We also recall Besicovitch’s covering theorem in the function field setup:
@@
[KT07, Example 2.1] Let $A$ be a bounded subset of $\mathcal{K}^{n}$. Then any covering by balls of $A$ has a countable subcover consisting of mutually disjoint balls.